OpenStudy (anonymous):

Lim [x] x to 0 --- ? x where [x] is the Greatest integer function

myininaya (myininaya):

what kinda function is that defined as, the one on top is the one i'm asking about?

OpenStudy (anonymous):

Greatest integer function

myininaya (myininaya):

is that one sided or actual limit?

OpenStudy (anonymous):

actual limit ? if it does not exist please give me the proper reason

OpenStudy (anonymous):

it must be zero because a zero divided by any number whatever small has to be zero.

myininaya (myininaya):

because greatest integer function of (x) =0 if 0<x<1 and greatest integer function of (x)=-1 if -1<x<0 Suppose we had \[\lim_{x \rightarrow 0^+}\frac{\lceil x \rceil}{x}=\lim_{x \rightarrow ^+}\frac{0}{x}=0\] but look from the other side and see what you

myininaya (myininaya):

you will see a cute surprise

OpenStudy (anonymous):

Thank you sir.

myininaya (myininaya):

\[\lim_{x \rightarrow 0^-}\frac{\lceil x \rceil}{x}\] what will the greatest integer function(x) approach if we are looking to the left of 0?

OpenStudy (anonymous):

inf

OpenStudy (anonymous):

right 1 am i ?

myininaya (myininaya):

perfect because we have -1/x

OpenStudy (anonymous):

so limit does not exist.!!

myininaya (myininaya):

and as x approaches 0 -1/x approaches inf the limit does not exist very good

myininaya (myininaya):

i think that all made sense to you? :)

OpenStudy (anonymous):

Yeah of course ..!!